OpenStudy (anonymous):

Need Help Given 21x^3-3x^2=0 , which values of x will satisfy the equation? (Hint: Factor the polynomial into two terms, and find the values of x that will make each term 0.) A. x = 0 and x = –7 B. x = 0 and x = 7 C. x=0 and x=-1/7 D. x=0 and x=1/7

OpenStudy (zehanz):

The two terms of this equation, \(21x^3\) and \(3x^2\), have common factors. If you factor each term first, you can see all the common factors: \(21x^3=3\cdot7\cdot x \cdot x\cdot x\) \(3x^2=3 \cdot x \cdot x\). You can factor out the common factors (which are 3, x and x, obviously). Now you have \(3x^2(7x-1)=0\). The equation now consists of two FACTORS and not terms btw. So solve \(3x^2=0\) and \(7x-1=0\) to find the right answer...